# REINFORCE++: An Efficient RLHF Algorithm with Robustness to Both Prompt and Reward Models

### v5

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) is crucial in aligning large language models (LLMs) with human values and preferences. While state-of-the-art applications like ChatGPT/GPT-4 commonly employ Proximal Policy Optimization (PPO), including a critic network introduces significant computational overhead. REINFORCE-based methods, such as REINFORCE Leave One-Out (RLOO), ReMax, and Group Relative Policy Optimization (GRPO), address this limitation by eliminating the critic network. However, these approaches face challenges in accurate advantage estimation. Specifically, they estimate advantages independently for responses to each prompt, which can lead to overfitting on more straightforward prompts and vulnerability to reward hacking. To address these challenges, we introduce REINFORCE++, a novel approach that removes the critic model while using the normalized reward of a batch as the baseline. Our empirical evaluation demonstrates that REINFORCE++ exhibits robust performance across various reward models without requiring prompt set truncation. Furthermore, it achieves superior generalization in RLHF and long chain-of-thought (CoT) settings compared to REINFORCE-based methods. The implementation is available at https://github.com/OpenRLHF/OpenRLHF.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is a key technique for aligning large language models (LLMs) with human values and preferences  (Vemprala et al., 2023; Achiam et al., 2023; Ouyang et al., 2022a; Shen & Zhang, 2024). Despite the emergence of non-RL alternatives like DPO  (Rafailov et al., 2023), state-of-the-art applications such as ChatGPT/GPT-4  (Vemprala et al., 2023; OpenAI, 2023), Claude  (Anthropic, 2023), and Gemini  (Team et al., 2023) continue to rely on RL algorithms, particularly PPO, for policy optimization. However, PPO  (Schulman et al., 2017) requires a critic network, introducing substantial computational overhead and memory demands that limit large model alignment in small-scale clusters. To address this, researchers have proposed various REINFORCE-based methods that eliminate the critic network, including ReMax  (Li et al., 2023), REINFORCE Leave One-Out (RLOO)  (Wu et al., 2024), and Group Relative Policy Optimization (GRPO)  (Shao et al., 2024). Furthermore, Deepseek-R1 demonstrates the effectiveness of the REINFORCE-based method in the long-form CoT setting  (Guo et al., 2025), which achieved state-of-the-art performance on challenging datasets using GRPO with rule-based rewards.
@@
Without the critic network, REINFORCE-based methods often struggle to estimate individual tokens’ advantages accurately. Various REINFORCE-baseline approaches have been proposed to address this limitation, but each has significant drawbacks. ReMax uses a greedy search to generate a response for each prompt and employs its reward as the baseline, inefficiently consuming a model response solely for baseline computation. RLOO and GRPO take a different approach by generating multiple responses per prompt: RLOO uses the mean reward of other responses as the baseline, while GRPO utilizes a normalized reward across all responses. These methods improve advantage estimation accuracy, but their practice of optimizing multiple responses per prompt intensifies the risk of reward hacking. Furthermore, these methods calculate reward baselines separately for each prompt, leading to overfitting and instability on specific training prompts during optimization. Therefore, these methods require carefully curating prompt sets particular to each task.
@@
To address these challenges, we propose REINFORCE++, a novel REINFORCE-based method that eliminates the critic model from PPO and uses the mean reward of a global batch as the baseline. This approach prevents overfitting to specific training prompts and demonstrates robustness across both Bradley-Terry and rule-based reward models. Notably, REINFORCE++ eliminates the need for prompt set truncation and achieves strong generalization performance in both RLHF and long CoT RL settings.
The remainder of this paper is organized as follows. In Section 2, we analyze the differences between various REINFORCE-based RLHF methods, demonstrating the ineffectiveness of prompt-specific reward baselines and the overfitting problems in RLOO and GRPO methods. Section 3 presents REINFORCE++, detailing its implementation and advantages. In Section 4, we evaluate REINFORCE++ through comprehensive experiments. We first compare REINFORCE++ with other RLHF methods using both Bradley-Terry and Rule-Based Reward Models, showing its superior or comparable performance. Then, under the long-form Chain-of-Thought (CoT) setting, we focus on comparing REINFORCE++ with GRPO, the current state-of-the-art REINFORCE-based method, due to computational constraints. Through experiments on a curated small set of carefully selected prompts, we highlight GRPO’s weakness in overfitting specific prompts. Our subsequent experiments on long-form CoT tasks demonstrate REINFORCE++’s superior performance on challenging test datasets, indicating better Out-of-Distribution (OOD) generalization. In Section 5, we conclude our method and future work directions.
@@
2 Background and Related Work
State-of-the-art applications such as ChatGPT/GPT-4, Claude, and Gemini utilize reinforcement learning algorithms, such as PPO, for policy optimization. In particular, PPO optimizes LLMs by maximizing the following surrogate objective:
$\mathcal{L}_{\text{PPO}}(\theta)=\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta_{\text{old}}}(O|q)}\left[\frac{1}{|o|}\sum_{t=1}^{|o|}\min\left(s_{t}(\theta)A_{t},\ \text{clip}(s_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right)\right]$ (1)
@@ 2 Background and Related Work
PPO is an actor-critic-based reinforcement learning (RL) algorithm where the critic model consumes significant training resources. Consequently, many researchers have proposed a series of REINFORCE-based methods, such as ReMax, RLOO, and GRPO, to avoid the computational overhead associated with the critic model while still obtaining relatively accurate token-wise advantage estimations. These methods design alternative techniques to calculate the baseline reward for each prompt as the advantage estimation.
Specifically, the advantage is computed to evaluate the performance of an action relative to a baseline in PPO. It starts by collecting samples of states, actions, rewards, and the following states. The advantage is then calculated using Generalized Advantage Estimation (GAE), which combines the temporal difference error $\delta_{q,o_{t}}=r_{t}+\gamma V(o_{t+1})-V(o_{t})$ over a series of time steps:
@@
where $\lambda$ is a parameter balancing bias and variance. To obtain an accurate advantage estimation and generalize to unseen tokens, PPO must use the critic model to learn the advantage function, obtaining an accurate advantage estimation and generalizing to unseen tokens.
@@
Notably, ReMax does not use this response to train the model.
Both RLOO and GRPO sample multiple responses for a prompt. RLOO adopts the average rewards of all other samples for the current prompt as a baseline, which is shown in Equation 6.
@@
However, given a diverse dataset, we consider that the baseline reward for each prompt is not essential in the RLHF setting. Although the baseline reward of each prompt can provide relatively accurate advantage estimation for each training prompt and help the model learn the response with the highest reward under a prompt, it also accelerates the reward hacking and overfitting problems. It is worth noting that RLHF has two key differences from traditional RL problems:
@@
Accordingly, both reward hacking and overfitting problems may degrade model performance. Specifically, optimizing multiple responses for a prompt in a batch using methods like RLOO and GRPO tends to overfit the best response under specific simple prompts, ultimately degrading the model’s generalization. Furthermore, optimizing multiple responses for a prompt within a single batch reduces the diversity of the model’s outputs, leading to a low diversity in token-level advantage distribution and causing overfitting on those tokens. In contrast, PPO does not suffer significantly from this issue, as the value network continues to be trained and retains the learned advantages, thereby supporting a more generalizable token-wise advantage.
@@ 3 Method
REINFORCE++ optimizes the PPO objective and employs the clipping strategy defined in Equation 1. To further reduce the variance of the gradient estimate, making the learning process more stable and efficient, we adopt the average reward of a global training batch as the baseline reward. Accordingly, the advantage of REINFORCE++ is the normalized reward-to-go in reinforcement learning:
@@
The gradient of the token-level KL penalty has been theoretically proven to be unbiased concerning the $k_{3}$ loss of GRPO in RLHF. Additionally, we normalize this advantage across the global batch for all prompts:
$A_{q,o_{t}}^{\text{norm}}=\frac{A_{q,o_{t}}-\text{mean}\left(A_{q,o_{t}}\right)}{\text{std}\left(A_{q,o_{t}}\right)},$ (10)
REINFORCE++ removes the prompt-specific reward baseline and does not optimize multiple responses from a single prompt within a training batch. Compared to PPO, REINFORCE++ eliminates the critic model and sets the discount factor to $1$. The detailed implementation of our algorithm is provided in Algorithm 1, with further algorithmic details discussed therein.
@@
4:     Sample a batch $\mathcal{D}_{b}$ from $\mathcal{D}$
@@
6:     Sample one output $o\sim\pi_{\text{old}}(\cdot\mid q)$ for each question $q\in\mathcal{D}_{b}$
@@
13:$\pi_{\theta}$
@@ 4 Experiments
The empirical evaluation of REINFORCE++ was conducted across diverse test scenarios to assess its performance comprehensively. In experiments using the Bradley-Terry Reward Model, we thoroughly compare REINFORCE++ against existing methods, including ReMax, RLOO, GRPO, and PPO. Due to computational constraints, we focused our comparisons between REINFORCE++ and GRPO, the current state-of-the-art REINFORCE-based method for experiments in the long-form CoT setting.
@@ 4.1 Performance with Outcome Reward Model
This study investigates reinforcement learning with human feedback (RLHF) in a general-domain setting. The experimental process began with an instruction-following policy model, fine-tuned on diverse datasets reflecting general-domain language tasks. The model was subsequently refined using reinforcement learning algorithms guided by a reward model. Specifically, a Bradley-Terry reward model was trained on human-generated preference data obtained through pairwise comparisons of model outputs. The reward model evaluates the policy model’s responses based on helpfulness, accuracy, coherence, and alignment with human intent and outputs an outcome reward value for each Bai et al. (2022).
@@
Score Length Per Token Score
@@
We use Chat-Arena-Hard Li et al. (2024) to evaluate our models. As illustrated in Table 1, GRPO achieves a slightly superior overall score of 46.8 compared to 46.7 for REINFORCE++. However, GRPO generates longer sequences of an average of 860 tokens, whereas REINFORCE++ produces shorter outputs of only 832 tokens. Consequently, when evaluating performance on a per-token basis, REINFORCE++ outperforms GRPO, achieving a higher per-token score of 0.0561 compared to GRPO’s 0.0544. The table suggests that REINFORCE++ provides more efficient outputs despite the preference bias of the evaluation model on the length.
@@
GSM8K MATH HumanEval MBPP Avg.

### v8

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) plays a crucial role in aligning large language models (LLMs) with human values and preferences. While state-of-the-art applications like ChatGPT or GPT-4 commonly employ Proximal Policy Optimization (PPO), the inclusion of a critic network introduces significant computational overhead. REINFORCE-based methods, such as REINFORCE Leave One-Out (RLOO), ReMax, and Group Relative Policy Optimization (GRPO), address this limitation by eliminating the critic network. However, these approaches face challenges in accurate advantage estimation. Specifically, they estimate advantages independently for responses to each prompt, which can lead to overfitting on simpler prompts and vulnerability to reward hacking and may be biased. To address these challenges, we introduce REINFORCE++, a novel approach that removes the critic model while using the global advantage normalization which is unbiased to improve the training stability. Our empirical evaluation demonstrates that REINFORCE++ exhibits robust performance across various reward models without requiring prompt set truncation. Furthermore, it achieves superior generalization in both RLHF and long chain-of-thought (CoT) settings compared to existing REINFORCE-based methods. The implementation is available at https://github.com/OpenRLHF/OpenRLHF.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is a key technique for aligning large language models (LLMs) with human values and preferences  (Vemprala et al., 2023; Achiam et al., 2023; Ouyang et al., 2022a; Shen & Zhang, 2024; Shen et al., 2025; Hu et al., 2024). Despite the emergence of non-RL alternatives like DPO  (Rafailov et al., 2023), state-of-the-art applications such as ChatGPT/GPT-4  (Vemprala et al., 2023; OpenAI, 2023), Claude  (Anthropic, 2023), and Gemini  (Team et al., 2023) continue to rely on RL algorithms, particularly PPO, for policy optimization. However, PPO  (Schulman et al., 2017) requires a critic network, introducing substantial computational overhead and memory demands that limit large model alignment in small-scale clusters. To address this, researchers have proposed various REINFORCE-based methods that eliminate the critic network, including ReMax  (Li et al., 2023), REINFORCE Leave One-Out (RLOO)  (Ahmadian et al., 2024), and Group Relative Policy Optimization (GRPO)  (Shao et al., 2024). Furthermore, Deepseek-R1 demonstrates the effectiveness of the REINFORCE-based method in the long-form CoT setting  (Guo et al., 2025; Seed et al., 2025), which achieved state-of-the-art performance on challenging datasets using GRPO with rule-based rewards.
@@
Without the critic network, REINFORCE-based methods often struggle to estimate individual tokens’ advantages accurately. Various REINFORCE-baseline approaches have been proposed to address this limitation, among which are those that introduce prompt-level baselines; yet each still has significant drawbacks. ReMax uses a greedy search to generate a response for each prompt and employs its reward as the baseline, inefficiently consuming a model response solely for baseline computation. RLOO and GRPO take a different approach by generating multiple responses per prompt: RLOO uses the mean reward of other responses as the baseline, while GRPO employs a normalized reward across all responses, this serves as a biased estimator of the advantage (see Appendix A). These methods improve advantage estimation accuracy, but their practice of optimizing multiple responses per prompt intensifies the risk of reward hacking. Furthermore, these methods calculate reward baselines separately for each prompt, leading to overfitting and instability on specific training prompts during optimization. Therefore, these methods require carefully curating prompt sets particular to each task.
@@
To address these challenges, we propose REINFORCE++, a novel REINFORCE-based method that eliminates the critic model from PPO and uses the global advantage normalization. This approach is unbiased and prevents overfitting to specific training prompts and demonstrates robustness across both Bradley-Terry and rule-based reward models. Notably, REINFORCE++ eliminates the need for prompt set truncation and achieves strong generalization performance in both RLHF and long CoT RL settings.
In summary, our contributions are as follows:
  • We analyze the limitations of existing REINFORCE-based RLHF methods, revealing that prompt-specific reward baselines are ineffective and identifying overfitting issues in RLOO and GRPO.
  • We propose REINFORCE++, a novel REINFORCE-based RLHF method, and detail its implementation, highlighting its advantages in addressing the identified limitations.
  • Through comprehensive experiments using both Bradley-Terry and Rule-Based Reward Models, we demonstrate that REINFORCE++ achieves superior or comparable performance compared to other RLHF methods.
  • In long-form Chain-of-Thought (CoT) settings, we show that GRPO suffers from overfitting to specific prompts, while REINFORCE++ exhibits better Out-of-Distribution (OOD) generalization, especially on challenging test datasets.
2 Background and Related Work
@@
State-of-the-art applications such as ChatGPT or GPT-4, Claude, and Gemini utilize reinforcement learning algorithms, such as PPO, for policy optimization. In particular, PPO optimizes LLMs by maximizing the following surrogate objective:
$\mathcal{L}_{\text{PPO}}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\ o\sim\pi_{\theta_{\text{old}}}(O|q)}\Bigg{[}\frac{1}{|o|}\sum_{t=1}^{|o|}\min\Big{(}s_{t}(\theta)A_{t},$
$\text{clip}(s_{t}(\theta),\ 1-\epsilon,\ 1+\epsilon)A_{t}\Big{)}\Bigg{]}$ (1)
@@
PPO is an actor-critic-based reinforcement learning algorithm where the critic model consumes significant training resources. Consequently, many researchers have proposed a series of REINFORCE-based methods, such as ReMax, RLOO, and GRPO, to avoid the computational overhead associated with the critic model while still obtaining relatively accurate token-wise advantage estimations. These methods design alternative techniques to calculate the baseline reward for each prompt as the advantage estimation.
Specifically, PPO computes the advantage to assess how an action’s return compares to the immediately preceding state value. It starts by collecting samples of states, actions, rewards, and the following states. The advantage is then calculated using Generalized Advantage Estimation (GAE), which combines the temporal difference error $\delta_{q,o_{t}}=r_{t}+\gamma V(o_{t+1})-V(o_{t})$ over a series of time steps:
@@
where $\lambda$ is a parameter balancing bias and variance. To obtain an accurate advantage estimation and generalize to unseen tokens, PPO use the critic model to learn the advantage function, obtaining an accurate advantage estimation and generalizing to unseen tokens.
@@
Notably, ReMax does not use this response to train the model. In addition, both RLOO and GRPO sample multiple responses for a prompt. RLOO adopts the average rewards of all other samples for the current prompt as a baseline, which is shown in Equation 6.
@@
For GRPO, which has been widely adopted in various works, we demonstrate that its advantage estimation is biased (see Appendix A for details). Furthermore, when working with a diverse dataset, we argue that the per-prompt baseline reward is not essential in the RLHF framework. While such a baseline can yield relatively accurate advantage estimates for each training prompt—thereby helping the model learn to generate responses with the highest reward under a given prompt—it also exacerbates reward hacking and overfitting issues. It is worth noting that RLHF has two key differences from traditional RL problems:
@@
Accordingly, both reward hacking and overfitting problems may degrade model performance. Specifically, optimizing multiple responses for a prompt in a batch using methods like RLOO and GRPO tends to overfit the best response under specific simple prompts, ultimately degrading the model’s generalization. Furthermore, when optimizing multiple responses for a single prompt within a training batch, the diversity of the model’s outputs during training is reduced. This leads to low diversity in the token-level advantage distribution, ultimately causing overfitting on those tokens. In contrast, PPO does not suffer significantly from this issue, as the value network continues to be trained and retains the learned advantages, thereby supporting a more generalizable token-wise advantage.
@@ 3 Method
REINFORCE++ still optimizes the PPO objective and employs the clipping strategy as defined in Equation 1. To further reduce the variance of the gradient estimate, making the learning process more stable and efficient, we adopt the average reward of a global training batch as the baseline reward. Accordingly, the advantage of REINFORCE++ is the normalized reward-to-go in reinforcement learning:
@@
Notably, Reinforce++ algorithm employs the KL-based $k_{1}$ loss. This choice is motivated by the fact that the GRPO algorithm, which relies on the KL-based $k_{3}$ loss, suffers from bias in its gradient estimation (For a detailed analysis, refer to Appendix KL Penalty Design.). Additionally, we normalize this advantage across the global batch for all prompts:
$A_{q,o_{t}}^{\text{norm}}=\frac{A_{q,o_{t}}-\text{mean}\left(A_{q,o_{t}}\mid A_{q,o_{t}}\in\mathcal{D}_{\text{batch}}\right)}{\text{std}\left(A_{q,o_{t}}\mid A_{q,o_{t}}\in\mathcal{D}_{\text{batch}}\right)}$ (10)
The global advantage normalization helps avoid training instability caused by excessively large advantage values (Andrychowicz et al., 2020). Since the global batch size is typically large, we can consider the mean and variance as constants without introducing bias into the policy gradient estimation. For further analysis, see the Appendix A.2. It is worth noting that, compared to PPO, REINFORCE++ essentially eliminates the critic model and sets the GAE discount factor to 111. The detailed implementation of our algorithm is provided in Algorithm 1, with further algorithmic details discussed in Appendix Algorithm Details.
@@
4:  Sample a batch $\mathcal{D}_{batch}$ from $\mathcal{D}$
@@
6:  Sample one output $o\sim\pi_{\text{old}}(\cdot\mid q)$ for each question $q\in\mathcal{D}_{batch}$
@@
[…]
3.1 REINFORCE++-Baseline
Recent research (Yue et al., 2025) showed that training models using multiple generated responses per prompt in reasoning tasks can further improve performance. Accordingly, we introduce a variant called REINFORCE++-Baseline, which integrates REINFORCE++ with multiple-response generation. Specifically, for each prompt we sample multiple responses and use their average reward as a baseline to reshape the rewards, therefore, under the RLVR setting, the reward function is not very sensitive whether it is 0/1 or -1/1. The advantage value for each prompt-response pair is then defined as:
$A_{q,o_{t}}=R_{q,o_{t}}-\text{mean}_{group}(R_{q,o_{t}})$ (11)
$A_{q,o_{t}}^{norm}=\frac{A_{q,o_{t}}-\text{mean}_{batch}(A_{q,o_{t}})}{\text{std}_{batch}(A_{q,o_{t}})}$ (12)
where $group$ denotes the generated responses corresponding to the same prompt. This baseline calculation is similar to the approach taken by GRPO; However, instead of using the local standard deviation in GRPO, we adopt the global batch normalization from REINFORCE++ to enhance training stability. For the REINFORCE++-baseline, we adopt the k2 KL estimator rather than k3 in GRPO, as k2 provides an unbiased estimate. See Appendix KL Penalty Design for more details.
3.2 Relationship with PPO
REINFORCE++ shares certain similarities with PPO in its formulation. Specifically, when PPO adopts Generalized Advantage Estimation (GAE) parameters with $\lambda=1$ and discount factor $\gamma=1$, REINFORCE++ reduces to PPO without the critic network and additionally employs global batch normalization as the baseline. Mathematically, this relationship is expressed as follows:
where $V(s_{t})$ denotes the estimated value function at state $s_{t}$. By removing the critic network, the REINFORCE++ algorithm effectively eliminates the $V(s_{t})$ term, while the introduction of global batch normalization further stabilizes training.
@@ 4 Experiments
The empirical evaluation of REINFORCE++ was conducted across diverse test scenarios to comprehensively assess its performance. In the context of RLHF, experiments utilizing the Bradley-Terry Reward Model were designed to thoroughly compare REINFORCE++ against existing methods, including ReMax, RLOO, GRPO, and PPO. For RLVR settings—specifically in long-form CoT experiments—our comparisons focused on REINFORCE++ versus GRPO, the current state-of-the-art REINFORCE-based method in this domain, due to computational constraints. We utilize binary (-1/1) rewards as training signals for the long-form CoT experiments. Additionally, we provide comparative experiments for our method under complex multi-turn tool call scenarios to demonstrate strong performance in agentic environments. This multi-faceted approach ensures a robust evaluation of REINFORCE++ and its variant across various reinforcement learning paradigms and application contexts. We outline the recommended best practices for our methods in Appendix D.
@@ 4.1 Performance with Outcome Reward Model
This study investigates RLHF in a general-domain setting. The experimental process began with an instruction-following policy model, fine-tuned on diverse datasets reflecting general-domain language tasks. The model was subsequently refined using reinforcement learning algorithms guided by a reward model. Specifically, a Bradley-Terry reward model was trained on human-generated preference data obtained through pairwise comparisons of model outputs. The reward model evaluates the policy model’s responses based on helpfulness, accuracy, coherence, and alignment with human intent and outputs an outcome reward value for each  (Bai et al., 2022).
@@
Score Length Per Token
@@
We use Chat-Arena-Hard (Li et al., 2024) to evaluate our models. As illustrated in Table 1, GRPO achieves a slightly superior overall score of 46.8 compared to 46.7 for REINFORCE++. However, GRPO generates longer sequences of an average of 860 tokens, whereas REINFORCE++ produces shorter outputs of only 832 tokens. Consequently, when evaluating performance on a per-token basis, REINFORCE++ outperforms GRPO, achieving a higher per-token score of 0.0561 compared to GRPO’s 0.0544. The table suggests that REINFORCE++ provides more efficient outputs despite the preference bias of the evaluation model on the length (Dubois et al., 2024).
@@
GSM8K MATH HumanRval MBPP Avg.
Base Model (before training) 95.83 68.80 82.71 82.2 82.39
@@ 4.2.3 RL from Zero Setting
4.2.4 Agent RL Setting: Tool-Integrated Reasoning and Agent Experiments
Experimental Setup
Our experiments are conducted within the framework established by Mai et al. (2025), which features a zero-shot agent environment designed for large language models to tackle mathematical problems. This environment allows the models to invoke external tools, such as Python execution, without the need for supervised demonstrations of tool usage.
To ensure a fair comparison, we adhere to the training and evaluation protocols of the ZeroTIR environment. The backbone model utilized in our experiments is Qwen 2.5 Base 7B, which is initialized from pre-trained weights to maintain consistency across trials. The training process is facilitated by OpenRLHF Hu et al. (2025a), utilizing datasets from ORZ Hu et al. (2025b) and DAPO Yu et al. (2025).
We assess the performance of the models on several established mathematical benchmarks, including AIME 2024 Jia , AIME 2025 math ai , HMMT FEB 2024/2025, and CMIMC. The evaluation metric employed is the average@32, consistent with the methodology outlined in Mai et al. (2025).
Experimental Result
The results of our experiments, summarized in Table 5, indicate that the REINFORCE++-baseline algorithm consistently outperforms both GRPO and PPO across all benchmarks, achieving the highest average accuracy of 24.10. In comparison, GRPO achieved an average accuracy of 22.58, while PPO lagged behind with an average accuracy of 21.85. This demonstrates the effectiveness of global batch advantage normalization in complex scenarios. For detailed scores across individual benchmarks, please refer to Appendix Tool-Integrated Reasoning and Agent Experiments.
5 Limitations
While REINFORCE++ demonstrates strong performance across diverse practical scenarios, it is important to contextualize its limitations with a balanced perspective on its strengths.
In iid (independent and identically distributed) settings, REINFORCE++ performs comparably to GRPO—often with improved stability. However, it does not clearly outperform GRPO in these structured environments, where GRPO’s mature optimization remains competitive. This reflects REINFORCE++’s intent as a stable alternative, not a dramatic performance leap, in iid contexts.
A more notable constraint arises from the absence of a critic model, which limits its performance ceiling relative to methods like PPO. PPO refines advantage estimates via a value function, while REINFORCE++ relies solely on reward-to-go. This simplicity aids stability but hinders precision in tasks with sparse or delayed rewards.
Furthermore, due to compute limitations, we cannot conclusively assess REINFORCE++ at larger training scales (e.g., 3k+ steps or 8k+ long chains of thought), where its final behavior remains uncertain. For example, Magistral (Rastogi et al., 2025) trained with a method quiet similar to REINFORCE++-baseline achieve stronger results under extended regimes, suggesting potential headroom for improvement.
Finally, REINFORCE++ builds on the REINFORCE family with baseline normalization. While it delivers strong empirical gains (especially in out-of-distribution (OOD) scenarios), it does not offer foundational theoretical innovations. Its contributions lie in practical refinements targeting known limitations of earlier variants.
These observations underscore that REINFORCE++ excels as a robust, stable solution—particularly in OOD contexts—while its role in iid or critic-heavy tasks should be weighed against specific needs.